I restarted c1scy, c1rfm (so both sender and receiver models were cycled) and power-cycled the c1iscey and c1sus machines. The TRY PD is certainly seeing light - it is just not getting piped over to c1rfm. dmesg doesn't give any clues. I'm out of ideas.
P.S. The new reality seems to be that getting ITMY stuck in the event of a c1susaux reboot is inevitable. As is the practise for ITMX, I tried slowly ramping the PIT and YAW biases to 0 slowly - but in the process of ramping YAW to 0, the optic got stuck. I am ramping in steps of 0.1 (in units of the PIT/YAW sliders, waiting ~3 seconds between steps), I guess I can try ramping even more slowly.
Update: I power cycled the physical RFM switch. This necessitated reboot of all vertex FEs. But seems like things are back to normal now...
Note: to unstick ITMY, seems like the best approach is:
The pressure is still 2e-4 torr according to CC1 so I thought I'd give ASS debugging a go tonight. But the arm transmission signal isn't coming through to the LSC model from the end PDs - so a resurfacing of this problem. Rebooting the sender model, c1scy, did not fix the problem. Moreover, c1susaux is dead. The last time I rebooted it, ITMY got stuck so I'm not going to attempt a revival tonight.